# Nonlinear Ringdowns as Sources and Detectors of Quantum Gravitational Waves

### v1

@@ Abstract
I Introduction
Recent years have seen an increased interest in experimental proposals aimed at detecting phenomena that require a unified view of quantum mechanics and gravity for their explanation [1, 2, 3, 4, 5, 6, 7, 8, 9]. While a complete quantum theory of gravity remains elusive [10, 11, 12], the effective field theory quantization of general relativity yields low-energy predictions that could in principle be experimentally verified. For instance, quantum gravitational waves (GWs) can occupy states with no counterpart in classical general relativity, such as the vacuum, single-graviton states, and squeezed or thermal states [13]. These states, some of which can have a macroscopic mean number of gravitons [14], can induce noise in interferometers [15, 13, 14, 16, 17], resonant bar detectors [18] and gravitational decoherence [19, 20, 21]. Detecting such effects would provide strong evidence for the quantum nature of the gravitational field. However, as emphasized in [22, 23], the detection of gravitational noise or graviton clicks would not rule out classical models of gravitational radiation with the same level of rigor as established by non-classicality witnesses in quantum optics [24]. To demonstrate the quantum nature of gravitational radiation on par with its electromagnetic counterpart, it is necessary to understand which non-classical GW states can be generated in nature and how they can be measured.
In quantum optics, non-classical states are generated by nonlinear processes [25, 26]. Since gravity can be strongly nonlinear [27], a natural question emerges: can the nonlinearities of Einstein’s theory also give rise to non-classicality? Recent numerical simulations have confirmed the presence of nonlinear effects in black hole (BH) ringdowns [28, 29]; see also [30, 31, 32] for earlier works. Here, we propose that nonlinearities during BH ringdowns may serve as a mechanism for generating GW states differing significantly from coherent states. BH mergers could therefore source non-classical gravitational radiation.
@@ I Introduction
Readers convinced that quantum properties of GWs cannot be detected at interferometers, or those only interested in GW nonlinear interactions, can skip Sec. V.
II Second harmonics in ringdowns
@@
$\mathcal{R}\equiv\frac{A_{(4,4)}^{(2,2,0)\times(2,2,0)}}{A_{(2,2,0)}^{2}},$ (1)
@@ III Hamiltonian
$H_{I}=\kappa a^{2}b^{\dagger}+\kappa^{*}a^{\dagger 2}b,$ (2)
@@
[…]
@@
After a suitable projection, the contributions from each mode can be isolated  [^1] As discussed for example in [85, 86]. Quantization then leads to annihilation and creation operators for the frequency modes of $\Psi^{(1,2)}$ [57, 58, 59]. We denote the annihilation operators for the $\omega$ and $2\omega$ modes as $a,b$. The source term $\mathcal{S}[\Psi^{(1)}]$ couples the eqs. of motion for $a$ and $b$, and since it depends quadratically on $\Psi^{(1)}$ and its first derivatives, these coupled eqs. must be derivable from a Hamiltonian containing terms at most cubic in the set $\{a,a^{\dagger},b,b^{\dagger}\}$  [^2] In fact, the quadratic dependence of the source term on $\Psi^{(1)}$ implies the Hamiltonian must be linear in $b$ and quadratic in $a$, since mode $b$ is ‘forced’ by quadratic functions of $\Psi^{(1)}$ and its derivatives, which yield quadratic expressions in $a$ and $a^{\dagger}$.
@@
The above procedure for deriving the interaction Hamiltonian for gravitational SHG was recently applied in a simplified setting in [61], but the same conclusions follow from the quadratic dependence of $\mathcal{S}$ on $\Psi^{(1)}$ and considerations of (approximate) energy conservation [62].
@@
Graviton number, or energy conservation requires that the Hamiltonian we are seeking commutes with the operator $J=a^{\dagger}a+2b^{\dagger}b$ . This is approximatelly valid for weakly damped modes, as well as for interaction times $t\ll\gamma^{-1}$. Writing down a sum of all possible terms up to third order in the modes’ creation and annihilation operators [as imposed by the form of (4) and source term], the conservation condition $[H,J]=0$ leads uniquely to (2) up to the coupling constant and a free field Hamiltonian. See the Appendix for details. In the gravitational context, the coupling constant $\kappa$ depends on spatial mode overlap integrals of the interacting GWs [63]. For simplicity, we assume $\kappa$ real.
@@
From the form of the interaction Hamiltonian, we can gain intuition on the quadratic dependence of the nonlinear signal on the amplitude of the parent mode. When mode $a$ is populated by a strong coherent state, we can effectively replace the operator $a$ by the corresponding c-number amplitude $\alpha_{(2,2,0)}\propto A_{(2,2,0)}$. This is known in quantum optics as the undepleted pump approximation [64]. Under this assumption, the interaction Hamiltonian becomes $H_{I}\sim\kappa\alpha_{(2,2,0)}^{2}b^{\dagger}+\mathrm{h.c.}$, which acts as a generator of displacements for mode $b$. The resulting evolution produces an approximate coherent state $|\alpha_{(4,4)}^{(2,2,0)\times(2,2,0)}\rangle\approx e^{-iH_{I}t}|0\rangle$ in $b$, with amplitude $\alpha_{(4,4)}^{(2,2,0)\times(2,2,0)}\propto A_{(4,4)}^{(2,2,0)\times(2,2,0)}=\mathcal{R}A_{(2,2,0)}^{2}$, and $\kappa\propto\mathcal{R}$.
@@
This coherent state picture of the nonlinear mode is a crude approximation, valid only in the undepleted pump approximation. It breaks down near $\mathcal{R}\sim 1$, when strong coupling leads to significant energy exchange between the parent and nonlinear modes. See the discussion on energy exchange below.
IV Equations of motion
@@
$\dot{\alpha}$ $=$ $-2i\kappa\beta\alpha^{\dagger}$ (5)
$\dot{\beta}$ $=$ $-i\kappa\alpha^{2}$ (6)
which are nonlinear coupled differential equations. We refer to the Appendix for the inclusion of damping terms. In quantum optics, it is known that these eqs. lead to squeezing and sub-Poissonian statistics of the parent and nonlinear modes [65, 66], which are nonclassical properties [25, 67]. Moreover, input parent coherent states undergoing SHG can also generate entanglement between the parent and nonlinear modes in the strong coupling regime, provided dissipation is not too large [68, 69, 70, 71, 72]. Hence, if GWs are quantized, nonlinear ringdowns generate non-classical gravitational radiation.
V Signatures on interferometers
Experimentally determining the quantum nature of gravity at the level of rigor of quantum optics, either for single mode or entangled states, requires the violation of a non-classicality witness, e.g. a Cauchy-Schwarz inequality [24] or entanglement witness [73]. We now investigate whether this could in principle be done by monitoring the quantum optical state of an interferometric detector interacting with a non-classical GW? The short answer is yes in principle, although as we will see, in practice direct measurements are extremely challenging.
Following [13, 74, 75], we consider an optical cavity as a model interferometric detector. Starting from the linearised Einstein’s equations in the Minkowski background [76], one can show that the interaction between the cavity field and GW modes consists of a dispersive optomechanical coupling in which the GWs formally assume the role of mechanical oscillators [13, 14]. We refer to this as the optogravitational interaction. As usual, we will work with a discrete set of GW modes in a cubic box of volume $V$ and take the continuum limit $V\rightarrow\infty$ when necessary. In order to encompass the case of bipartite entangled states, we consider two GW modes, but note this can be extended to the case of single-mode non-classical states or multi-partite entangled states in a straightforward manner.
@@ V Signatures on interferometers
$H_{O}=\omega_{c}c^{\dagger}c+\omega_{a}a^{\dagger}a+\omega_{b}b^{\dagger}b-c^{\dagger}c\left(g_{a}Q_{a}+g_{b}Q_{b}\right)\ ,$ (7)
where $c(c^{\dagger})$ represents the annihilation (creation) operator for the cavity, $\omega_{c},\omega_{a,b}$ are the cavity and GWs’ frequencies, respectively and $Q_{a}=a+a^{\dagger},Q_{b}=b+b^{\dagger}$ are the amplitude quadratures for the GW modes. We also define the GW’s phase quadratures according to $P_{a}=i(a^{\dagger}-a)$ and $P_{b}=i(b^{\dagger}-b)$. The couplings $g_{i}$ are given by
$g_{i}\equiv\omega_{i}q_{i}=\frac{\omega_{c}}{4}\sqrt{\frac{8\pi G}{\omega_{i}V}}\ ,$ (8)
@@
$q_{i}\sim(\omega_{c}/E_{\rm pl})\ll 1\ ,$ (9)
which means GW detectors are weakly coupled optomechanical systems [13], i.e. GW detectors are weakly gravitating systems [34].
The Hamiltonian (7) can be exponentiated to give the time-evolution operator, which in the interaction picture reads [77]
$U(t)=e^{-iB(t)(c^{\dagger}c)^{2}}e^{q_{a}c^{\dagger}c(\eta_{a}a^{\dagger}-\eta_{a}^{*}a)}e^{q_{b}c^{\dagger}c(\eta_{b}b^{\dagger}-\eta_{b}^{*}b)}$ (10)
@@
$\eta_{k}$ $=$ $(1-e^{-i\omega_{k}t})$ (11)
$B(t)$ $=$ $\sum_{k=a,b}q_{k}^{2}\left(\omega_{k}t-\sin\omega_{k}t\right)\ .$ (12)
We refer to [13] and the Appendix for details.
@@
$\rho(0)=\left(\sum_{n,m}a_{n}a^{*}_{m}|n\rangle\langle m|\right)\otimes\sigma_{ab}\ ,$ (13)
where $a_{n}$ are known complex coefficients defining the initial optical density matrix and $\sigma_{ab}$ is the GWs’ initial state. Using the time-evolution operator (10), we can compute the reduced optical density matrix $\rho_{c}(t)=\mathrm{Tr}_{ab}\left(U\rho_{abc}U^{\dagger}\right)$. In the photon-number basis we find,
$[\rho_{c}(t)]_{nm}\equiv\langle n|\rho_{c}(t)|m\rangle=a_{n}a^{*}_{m}e^{-i\delta B(t)}e^{-i\omega_{c}\Delta t}\ \mathcal{C}_{\Delta}(t)\ ,$ (14)
@@
$\mathcal{C}_{\Delta}(t)=\mathrm{Tr}\left(\sigma_{ab}e^{\Delta q_{a}(\eta_{a}^{*}a^{\dagger}-\eta_{a}a))}e^{\Delta q_{b}(\eta_{b}^{*}b^{\dagger}-\eta_{b}b))}\right)\ .$ (15)
@@
$\mathcal{C}_{\Delta}(t)\approx 1$ - $iq_{a}\Delta\left(\sin\omega_{a}t\langle Q_{a}\rangle+(1-\cos\omega_{a}t)\langle P_{a}\rangle\right)$ (16)
- $\frac{q_{a}^{2}\Delta^{2}}{2}\left(\sin^{2}\omega_{a}t\langle Q_{a}^{2}\rangle+(1-\cos\omega_{a}t)^{2}\langle P_{a}^{2}\rangle+\sin\omega_{a}t(1-\cos\omega_{a}t)\bigl{<}Q_{a}P_{a}+P_{a}Q_{a}\bigr{>}\right)$
$+$ $(a\rightarrow b)$
- $q_{a}q_{b}\Delta^{2}\big{(}\sin\omega_{a}t\sin\omega_{b}t\langle Q_{a}Q_{b}\rangle+\sin\omega_{a}t(1-\cos\omega_{b}t)\langle Q_{a}P_{b}\rangle$
$\ +\sin\omega_{b}t(1-\cos\omega_{a}t)\langle P_{a}Q_{b}\rangle+(1-\cos\omega_{a}t)(1-\cos\omega_{b}t)\langle P_{a}P_{b}\rangle\big{)}\ ,$
@@
$\mathcal{D}=\mathrm{Var}(Q_{a}+Q_{b})+\mathrm{Var}(P_{a}-P_{b})\ ,$ (17)
@@
$\langle\gamma(t)\rangle=\mathcal{C}(t)\langle\gamma(0)\rangle+(\mathrm{terms\ independent\ of}\ \sigma_{ab})$ (18)
@@
$\langle X(t)\rangle$ $\approx$ $\mathrm{Re}(\mathcal{C}(t))\ \langle X(0)\rangle-\mathrm{Im}(\mathcal{C}(t))\ \langle Y(0)\rangle$ (19)
$\langle Y(t)\rangle$ $\approx$ $\mathrm{Im}(\mathcal{C}(t))\ \langle X(0)\rangle+\mathrm{Re}(\mathcal{C}(t))\ \langle Y(0)\rangle$ (20)
@@
where we have ignored terms independent of $\sigma_{ab}$ in (18). For small couplings $q_{a,b}$, this corresponds to a rotation in the optical phase space. To see that, consider for simplicity an initial GW state of the form $\sigma_{ab}=\rho_{a}\otimes|0\rangle\langle 0|_{b}$ and an initial optical coherent state with $\langle Y_{c}\rangle=0$. To linear order in $q_{a}$,
$\langle X(t)\rangle$ $\approx$ $\langle X(0)\rangle$ (21)
$\langle Y(t)\rangle$ $\approx$ $-q_{a}\left(\sin\omega_{a}t\langle Q_{a}\rangle-\cos\omega_{a}t\langle P_{a}\rangle\right)\langle X(0)\rangle$
@@
where we have neglected constant terms and vacuum fluctuations from mode $b$; see the Appendix for a discussion. The mean phase space position of the optical field oscillates in proportion to the GW field quadratures at the frequency of the wave. The strength of the oscillations is directly proportional to the coupling constant $q_{a}$ and the cavity quadrature $\langle X(0)\rangle$.
@@
This is precisely what is observed in interferometric detectors [13]. We conclude that, to linear order in the coupling $q_{a}$, interferometric detectors only have access to the mean field quadratures of the GW. Looking at (16), we see that second moments of the GW appear at $\mathcal{O}(q_{a}^{2})$ in $\mathcal{C}(t)$, and hence will only affect the optical quadratures at quadratic orders in the coupling.
@@
Given the smallness of the coupling constant (9), this shows that detecting second moments of the GWs would require monitoring the optical cavity with precision vastly greater than achieved in today’s detectors and very likely out of reach from conceivable experiments. Measuring classical gravitational waves, which is a linear effect in $q_{a}$, is at the limit of current technology; imagine detecting effects which are quadratic in $q_{a}$. This is a reflection of the weak coupling between matter to gravity, which makes GW detectors weakly gravitating systems. To detected quantum features of the gravitational field, we need strongly gravitating quantum systems.
VI Ringdown nonlinearities as detectors

### v2

@@ Abstract
1 Introduction
Recent years have seen an increased interest in experimental proposals aimed at detecting phenomena that require a unified view of quantum mechanics and gravity for their explanation [1, 2, 3, 4, 5, 6, 7, 8, 9]. While a complete quantum theory of gravity remains elusive [10, 11, 12], the effective field theory quantization of general relativity yields low-energy predictions that could in principle be experimentally verified. For instance, quantum gravitational waves (GWs) can occupy states with no counterpart in classical general relativity, such as the vacuum, single-graviton states, and squeezed or thermal states [13]. These states, some of which can have a macroscopic mean number of gravitons [14], can induce noise in interferometers [15, 13, 14, 16, 17], resonant bar detectors [18] and gravitational decoherence [19, 20, 21]. Detecting such effects would provide strong evidence for the quantum nature of the gravitational field. However, as emphasized in [22, 23], the detection of gravitational noise or graviton clicks would not rule out classical models of gravitational radiation with the same level of rigor as established by non-classicality witnesses in quantum optics [24]. To demonstrate the quantum nature of gravitational radiation on par with its electromagnetic counterpart, it is necessary to violate a non-classicality witness, for example either by detecting entanglement or sub-Poissonian statistics [25]. Determining whether this is possible will require understanding which non-classical GW states can be generated in nature and how they can be measured.
In quantum optics, non-classical states are generated by nonlinear processes [26, 27]. Since gravity can be strongly nonlinear [28], a natural question emerges: can the nonlinearities of Einstein’s theory also give rise to non-classicality? Recent numerical simulations have confirmed the presence of nonlinear effects in black hole (BH) ringdowns [29, 30]; see also [31, 32, 33] for earlier works. Here, we propose that nonlinearities during BH ringdowns may serve as a mechanism for generating GW states differing significantly from coherent states. BH mergers could therefore source non-classical gravitational radiation  [^1] Beyond effective quantization of general relativity and the standard model, new physics could also lead to the generation of non-classical GWs [34].
@@ 1 Introduction
Readers convinced that quantum properties of GWs cannot be detected at interferometers, or those only interested in GW nonlinear interactions, can skip section 5.
2 Second harmonics in ringdowns
@@
$\mathcal{R}\equiv\frac{A_{(4,4)}^{(2,2,0)\times(2,2,0)}}{A_{(2,2,0)}^{2}},$ (2.1)
@@ 3 Hamiltonian
$H_{I}=\kappa a^{2}b^{\dagger}+\kappa^{*}a^{\dagger 2}b,$ (3.2)
@@
The perturbation eq. (3.4) admits separation of variables: the angular part is expressed in tensor harmonics for Schwarzschild and spin-weighted spheroidal harmonics for Kerr, the radial part satisfies time-independent Schrödinger eqs. with appropriate boundary conditions [53], while the temporal part corresponds to a discrete set of damped harmonic oscillators characterized by the QNMs’ frequencies [54]. We focus on two modes with frequencies $\omega$ (component of $\Psi^{(1)}$) and $2\omega$ (component of $\Psi^{(2)}$). These correspond to the $\ell=m=2$ and $\ell=m=4$ modes in the case of nonlinear ringdowns. We now assume weak damping, $\gamma\equiv\Im\{\omega\}\ll\Re\{\omega\}$, which is a good approximation for rapidly rotating BHs, with zero-damping modes being the extreme case [55, 56, 57], and in the eikonal limit [58, 59].
@@
After a suitable “projection”, the contributions from each mode can be isolated  [^2] As discussed for example in [60, 61]. Quantization then leads to annihilation and creation operators for the frequency modes of $\Psi^{(1,2)}$ [62, 63, 64]. We denote the annihilation operators for the $\omega$ and $2\omega$ modes as $a,b$, respectively. The source term $\mathcal{S}[\Psi^{(1)}]$ couples the eqs. of motion for $a$ and $b$, and since it depends quadratically on $\Psi^{(1)}$ and its first derivatives, these coupled eqs. must be derivable from a Hamiltonian containing terms at most cubic in the set $\{a,a^{\dagger},b,b^{\dagger}\}$  [^3] In fact, the quadratic dependence of the source term on $\Psi^{(1)}$ implies the Hamiltonian must be linear in $b$ and quadratic in $a$, since mode $b$ is “forced” by quadratic functions of $\Psi^{(1)}$ and its derivatives, which yield quadratic expressions in $a$ and $a^{\dagger}$.
@@
The above procedure for deriving the interaction Hamiltonian for gravitational SHG was recently applied in a simplified setting in [65], but the same conclusions follow from the quadratic dependence of $\mathcal{S}$ on $\Psi^{(1)}$ and considerations of approximate energy conservation [66].
@@
Graviton number, or energy conservation requires that the Hamiltonian we are seeking commutes with the operator $J=a^{\dagger}a+2b^{\dagger}b$ . This is approximatelly valid for weakly damped modes, as well as for interaction times $t\ll\gamma^{-1}$. Writing down a sum of all possible terms up to third order in the modes’ creation and annihilation operators [as imposed by the form of (3.4) and source term], the conservation condition $[H,J]=0$ leads uniquely to (3.2) up to the coupling constant and a free field Hamiltonian. See A for details. In the gravitational context, the coupling constant $\kappa$ depends on spatial mode overlap integrals of the interacting GWs [67]. For simplicity, we assume $\kappa$ real.
@@
From the form of the interaction Hamiltonian, we can gain intuition on the quadratic dependence of the nonlinear signal on the amplitude of the parent mode. When mode $a$ is populated by a strong coherent state, we can effectively replace the operator $a$ by the corresponding c-number amplitude $\alpha_{\omega}\propto A_{(2,2,0)}$. This is known in quantum optics as the undepleted pump approximation [68]. Under this assumption, the interaction Hamiltonian becomes $H_{I}\sim\kappa\alpha_{\omega}^{2}\ b^{\dagger}+\mathrm{h.c.}$, which acts as a generator of displacements for mode $b$. The resulting evolution produces an approximate coherent state $|\alpha_{2\omega}\rangle\approx e^{-iH_{I}t}|0\rangle$ in $b$, where
$\alpha_{2\omega}\sim\kappa t\ \alpha_{\omega}^{2}$ (3.5)
and $\alpha_{2\omega}\propto A_{(4,4)}^{(2,2,0)\times(2,2,0)}=\mathcal{R}A_{(2,2,0)}^{2}$. We thus see that $\kappa\propto\mathcal{R}$. Later on, in section 6 we will use (3.5) to estimate properties of the GW states resulting from nonlinear interactions.
@@
This coherent state picture of the nonlinear mode is a crude approximation, valid only in the undepleted pump approximation. It breaks down near $\mathcal{R}\sim 1$, when strong coupling leads to significant energy exchange between the parent and nonlinear modes.
4 Equations of motion
@@
$\dot{\alpha}=-2i\kappa\beta\alpha^{\dagger}$ (4.6)
$\dot{\beta}=-i\kappa\alpha^{2}$ (4.7)
which are nonlinear coupled differential equations. We refer to B for the inclusion of damping terms. In quantum optics, it is known that these eqs. lead to squeezing and sub-Poissonian statistics of the parent and nonlinear modes [69, 70], which are nonclassical properties [26, 25]. Moreover, input parent coherent states undergoing SHG can also generate entanglement between the parent and nonlinear modes in the strong coupling regime, provided dissipation is not too large [71, 72, 73, 74, 75]. Hence, if GWs are quantized, nonlinear ringdowns generate non-classical gravitational radiation.
5 Signatures on interferometers
Experimentally determining the quantum nature of gravity at the level of rigor of quantum optics, either for single mode or entangled states, requires the violation of a non-classicality witness, e.g. a Cauchy-Schwarz inequality [24] or entanglement witness [76]. We now investigate whether this could in principle be done by monitoring the quantum optical state of an interferometric detector interacting with a non-classical GW. The short answer is yes in principle, although as we will see, in practice direct measurements are extremely challenging. This is consistent with [22, 23].
Following [13, 14], we consider an optical cavity as a model interferometric detector. Starting from the linearised Einstein’s equations in the Minkowski background [77], one can show that the interaction between the cavity field and GW modes consists of a dispersive optomechanical coupling in which the GWs formally assume the role of mechanical oscillators [13, 14]. We refer to this as the optogravitational interaction. As usual, we will work with a discrete set of GW modes in a cubic box of volume $V$ and take the continuum limit $V\rightarrow\infty$ when necessary. In order to encompass the case of bipartite entangled states, we consider two GW modes, but note this can be extended to the case of single-mode non-classical states or multi-partite entangled states in a straightforward manner.
@@ 5 Signatures on interferometers
$H_{O}=\omega_{c}c^{\dagger}c+\omega_{a}a^{\dagger}a+\omega_{b}b^{\dagger}b-c^{\dagger}c\left(g_{a}Q_{a}+g_{b}Q_{b}\right)\ ,$ (5.8)
where $c(c^{\dagger})$ represents the annihilation (creation) operator for the cavity; $\omega_{c},\omega_{a,b}$ are the cavity and GWs’ frequencies, respectively and $Q_{a}=a+a^{\dagger},Q_{b}=b+b^{\dagger}$ are the amplitude quadratures for the GW modes. We also define the GW’s phase quadratures according to $P_{a}=i(a^{\dagger}-a)$ and $P_{b}=i(b^{\dagger}-b)$. The couplings $g_{i}$ are given by
$g_{i}\equiv\omega_{i}q_{i}=\frac{\omega_{c}}{4}\sqrt{\frac{8\pi G}{\omega_{i}V}}\ ,$ (5.9)
@@
$q_{i}\sim(\omega_{c}/E_{\rm pl})\ll 1\ ,$ (5.10)
which hints that GW detectors are weakly coupled optomechanical systems [13], i.e. GW detectors are weakly gravitating systems [36].
The Hamiltonian (5.8) can be exponentiated to give the time-evolution operator, which in the interaction picture reads [78],
$U(t)=e^{-iB(t)(c^{\dagger}c)^{2}}e^{q_{a}c^{\dagger}c(\eta_{a}a^{\dagger}-\eta_{a}^{*}a)}e^{q_{b}c^{\dagger}c(\eta_{b}b^{\dagger}-\eta_{b}^{*}b)}$ (5.11)
@@
$\eta_{k}=(1-e^{-i\omega_{k}t})$ (5.12)
$B(t)=\sum_{k=a,b}q_{k}^{2}\left(\omega_{k}t-\sin\omega_{k}t\right)\ .$ (5.13)
We refer to [13] and C for details.
@@
$\rho(0)=\left(\sum_{n,m}a_{n}a^{*}_{m}|n\rangle\langle m|\right)\otimes\sigma_{ab}\ ,$ (5.14)
where $a_{n}$ are known complex coefficients defining the initial optical density matrix and $\sigma_{ab}$ is the GWs’ initial state. Using the time-evolution operator (5.11), we can compute the reduced optical density matrix $\rho_{c}(t)=\mathrm{Tr}_{ab}\left(U\rho_{abc}U^{\dagger}\right)$. In the photon-number basis we find,
$\rho_{c}(t)]_{nm}\equiv\langle n|\rho_{c}(t)|m\rangle=a_{n}a^{*}_{m}e^{-i\delta B(t)}e^{-i\omega_{c}\Delta t}\ \mathcal{C}_{\Delta}(t)\ ,$ (5.15)
@@
$\mathcal{C}_{\Delta}(t)=\mathrm{Tr}\left(\sigma_{ab}e^{\Delta q_{a}(\eta_{a}^{*}a^{\dagger}-\eta_{a}a))}e^{\Delta q_{b}(\eta_{b}^{*}b^{\dagger}-\eta_{b}b))}\right)\ .$ (5.16)
@@
$\mathcal{C}_{\Delta}(t)\approx 1-iq_{a}\Delta\left(\sin\omega_{a}t\langle Q_{a}\rangle+(1-\cos\omega_{a}t)\langle P_{a}\rangle\right)$
$\qquad-\frac{q_{a}^{2}\Delta^{2}}{2}\Bigl(\sin^{2}\omega_{a}t\langle Q_{a}^{2}\rangle+(1-\cos\omega_{a}t)^{2}\langle P_{a}^{2}\rangle$
[…]
$\qquad+(a\rightarrow b)$
$\qquad-q_{a}q_{b}\Delta^{2}\Bigl(\sin\omega_{a}t\sin\omega_{b}t\langle Q_{a}Q_{b}\rangle+\sin\omega_{a}t(1-\cos\omega_{b}t)\langle Q_{a}P_{b}\rangle$
$\qquad\qquad\qquad+\sin\omega_{b}t(1-\cos\omega_{a}t)\langle P_{a}Q_{b}\rangle+(1-\cos\omega_{a}t)(1-\cos\omega_{b}t)\langle P_{a}P_{b}\rangle\Bigr),$ (5.17)
@@
$\mathcal{D}=\mathrm{Var}(Q_{a}+Q_{b})+\mathrm{Var}(P_{a}-P_{b})\ ,$ (5.18)
@@
$\langle\gamma(t)\rangle=\mathcal{C}(t)\langle\gamma(0)\rangle+(\mathrm{terms\ independent\ of}\ \sigma_{ab})\ ,$ (5.19)
@@
$\langle X(t)\rangle$ $\approx$ $\mathrm{Re}(\mathcal{C}(t))\ \langle X(0)\rangle-\mathrm{Im}(\mathcal{C}(t))\ \langle Y(0)\rangle$ (5.20)
$\langle Y(t)\rangle$ $\approx$ $\mathrm{Im}(\mathcal{C}(t))\ \langle X(0)\rangle+\mathrm{Re}(\mathcal{C}(t))\ \langle Y(0)\rangle$ (5.21)
@@
where we have ignored terms independent of $\sigma_{ab}$ in (5.19). For small couplings $q_{a,b}$, this corresponds to a rotation in the optical phase space. To see that, consider for simplicity an initial GW state of the form $\sigma_{ab}=\rho_{a}\otimes|0\rangle\langle 0|_{b}$ and an initial optical coherent state with $\langle Y_{c}\rangle=0$. To linear order in $q_{a}$,
$\langle X(t)\rangle$ $\approx$ $\langle X(0)\rangle$ (5.22)
$\langle Y(t)\rangle$ $\approx$ $-q_{a}\left(\sin\omega_{a}t\langle Q_{a}\rangle-\cos\omega_{a}t\langle P_{a}\rangle\right)\langle X(0)\rangle$ (5.23)
@@
where we have neglected constant terms and vacuum fluctuations from mode $b$; see C for a discussion. The mean phase space position of the optical field oscillates in proportion to the GW field quadratures at the frequency of the wave. The strength of the oscillations is directly proportional to the coupling constant $q_{a}$ and the cavity quadrature $\langle X(0)\rangle$.
@@
This is precisely what is observed in interferometric detectors [13]. We conclude that, to linear order in the coupling $q_{a}$, interferometric detectors only have access to the mean field quadratures of the GW. Looking at (5.17), we see that second moments of the GW appear at $\mathcal{O}(q_{a}^{2})$ in $\mathcal{C}(t)$, and hence will only affect the optical quadratures at quadratic orders in the coupling.
@@
Given the smallness of the coupling constant (5.10), this shows that detecting second moments of the GWs would require monitoring the optical cavity with precision vastly greater than achieved in today’s detectors and very likely out of reach from conceivable experiments. Measuring classical gravitational waves, which is a linear effect in $q_{a}$, is at the limit of current technology; imagine detecting effects which are quadratic in $q_{a}$. This is a reflection of the weak coupling between matter to gravity, which makes GW detectors weakly gravitating systems. To detected quantum features of the gravitational field, we need strongly gravitating quantum systems.
6 Ringdown nonlinearities as detectors